OpenStudy (anonymous):

4. A right triangle has a side length that measures 4 m and a hypotenuse that measures 8.5 m. What is the measure of the other side of the triangle? (1 point) 7.5 m 8.1 m 9.4 m 9.8 m

OpenStudy (anonymous):

\[a^2+b^2=h^2\] \[b^2=h^2-a^2\] \[b=\sqrt{h^2-a^2}\]

OpenStudy (anonymous):

in your case \[b=\sqrt{8.5^2-4^2}\]

OpenStudy (anonymous):

i get \(7.5\) http://www.wolframalpha.com/input/?i=sqrt%288.5^2-4^2%29

OpenStudy (anonymous):

i got 4.5

OpenStudy (anonymous):

this is a 8:15:17 triangle in disguise each number was cut in half

OpenStudy (anonymous):

ohok i see..thanks :]]